OpenStudy (anonymous):

during 2000, two of the top-grossing concert tours were kiss and the dave matthews band. together the two visited 163 cities. kiss visited 77 more cities than the dave matthews band. how many cities did each group visited

Directrix (directrix):

K + M = 163 where K is the number of cities Kiss visited and M is the number of cities visited by Dave Matthews. K = M + 77 This comes from: kiss visited 77 more cities than the dave matthews band.

Directrix (directrix):

K + M = 163 K = M + 77 ----------- The task is to solve these two equations simultaneously. I recommend substitution. K + M = 163 Replace K by M + 77 (M + 77) + M = 163 Solve that for M. @Bweezy5

Directrix (directrix):

(M + 77) + M = 163 Solve that for M. @Bweezy5

OpenStudy (anonymous):

thank you!

Directrix (directrix):

You are welcome.